OpenStudy (anonymous):

Tiara wants to divide a square piece of paper into two equalivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure? A. 8.9 cm B. 20 cm C. 40 cm D. 28.3 cm

OpenStudy (ivettef365):

use Pythagorean Theorem 20^2 + 20^2 = c^2 800 = c^2 what is the sqrt800

OpenStudy (anonymous):

8x100?

OpenStudy (ivettef365):

use calculator \(\sqrt{800}\)

OpenStudy (anonymous):

10

OpenStudy (ivettef365):

or you can multiply 20 x \(\sqrt{2}\)

OpenStudy (anonymous):

i dont have a calculator that does that

OpenStudy (ivettef365):

http://www.picalc.com/

OpenStudy (anonymous):

i dont know how to work it how do i mak the check

OpenStudy (anonymous):

im sorry i havent done math in 10yrs

OpenStudy (ivettef365):

sqrt800 = 28.2842

OpenStudy (anonymous):

so the answer would be d

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thanks u